OpenStudy (anonymous):

Find the equation in general form of the line (5,-4) (-4,3)

OpenStudy (a_clan):

Ax+By=C

OpenStudy (anonymous):

find the slope of the line find the intercept put i the form y = mx +b done

OpenStudy (anonymous):

\[y = -(7 x)/9-1/9\]

OpenStudy (anonymous):

What if none of the answers are fractions?

OpenStudy (anonymous):

Then it would be: \[7 x+9 y+1 = 0\]

OpenStudy (anonymous):

7x+9y=-1 or 9x+7y=-1 or 5x-4y=0 or -4x+3y=0 which of these would be the answer?

OpenStudy (a_clan):

first

OpenStudy (anonymous):

why is it -1?

OpenStudy (a_clan):

7x+9y+1=0 convert it in Ax+By=C form: 7x+9y=-1

OpenStudy (anonymous):

Oh, ok! May I ask how you got the answer?

OpenStudy (anonymous):

Like the process

OpenStudy (a_clan):

AS duke already explained: find slope, m=(y2-y1)/(x2-x1) Put it in y=mx+c then put any one co-ordinate (x1,y1) or (x2,y2) in this equation, you will get c. Put back value of c in equation, you will get y=-(7x)/9 -1/9 Re-arrange this equation into standard form:Ax+ By=C you will have 7x+9y=-1

OpenStudy (anonymous):

when we try and find the slope, we come up with -7/9

OpenStudy (a_clan):

correct

OpenStudy (anonymous):

what numbers do we use to plug it in?